# MRFD: Multi-Region Fusion Decoding with Self-Consistency for Mitigating Hallucinations in LVLMs

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have shown strong performance across multimodal tasks. However, they often produce hallucinations—text that is inconsistent with visual input, due to the limited ability to verify information in different regions of the image. To address this, we propose Multi-Region Fusion Decoding (MRFD), a training-free decoding method that improves factual grounding by modeling inter-region consistency. MRFD identifies salient regions using cross-attention, generates initial responses for each, and computes reliability weights based on Jensen-Shannon Divergence (JSD) among the responses. These weights guide a consistency-aware fusion of per-region predictions, using region-aware prompts inspired by Chain-of-Thought reasoning. Experiments across multiple LVLMs and benchmarks show that MRFD significantly reduces hallucinations and improves response factuality without requiring model updates.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have emerged as powerful tools for multimodal understanding, achieving significant advances in image captioning, visual question answering, and visual reasoning  Alayrac et al. (2022); Liu et al. (2023b); Li et al. (2023a); Zhu et al. (2023). However, these models frequently generate hallucinated content, producing textual outputs inconsistent with the visual input. This phenomenon poses substantial challenges for reliability and practical deployment Ji et al. (2023); Liu et al. (2024). Hallucinations typically manifest as misidentified objects, fabricated attributes, or omitted visual information, reducing the factual alignment between image and text.
@@ Appendix F Efficiency Comparison
[…]
Method Avg. Lat. $\downarrow$ GPU Mem. $\downarrow$ Cs $\downarrow$
Regular 3.44 s 15778 MB 55.0
VCD 6.91 s 16634 MB 54.4
OPERA 24.70 s 22706 MB 52.6
Woodpecker 10.68 s 22199 MB 57.6
HALC 22.61 s 23084 MB 51.0
DeGF 13.89 s 19119 MB 48.8
Ours (MRFD) 10.21 s 16932 MB 34.6
Table 8: Efficiency comparison on InstructBLIP. Lower is better for all metrics.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) have shown strong performance across multimodal tasks. However, they often produce hallucinations—text that is inconsistent with visual input, due to the limited ability to verify information in different regions of the image. To address this, we propose Multi-Region Fusion Decoding (MRFD), a training-free decoding method that improves factual grounding by modeling inter-region consistency. MRFD identifies salient regions using cross-attention, generates initial responses for each, and computes reliability weights based on Jensen-Shannon Divergence (JSD) among the responses. These weights guide a consistency-aware fusion of per-region predictions, using region-aware prompts inspired by Chain-of-Thought reasoning. Experiments across multiple LVLMs and benchmarks show that MRFD significantly reduces hallucinations and improves response factuality without requiring model updates. Code is available at https://github.com/Haonan-Ge/MRFD.
MRFD: Multi-Region Fusion Decoding with Self-Consistency for Mitigating Hallucinations in LVLMs
Haonan Geλ   Yiwei Wangλthanks:   Corresponding author.   Ming-Hsuan Yangλ   Yujun Cai λ Department of Computer Science and Engineering, University of California at Merced The University of Queensland gehaonan82@gmail.comwangyw.evan@gmail.com https://mrfd1.github.io
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have emerged as powerful tools for multimodal understanding, achieving significant advances in image captioning, visual question answering, and visual reasoning  Alayrac et al. (2022); Liu et al. (2023b); Li et al. (2023a); Zhu et al. (2023); Xia et al. (2024); Li et al. (2025b, a). However, these models frequently generate hallucinated content, producing textual outputs inconsistent with the visual input. This phenomenon poses substantial challenges for reliability and practical deployment Ji et al. (2023); Liu et al. (2024). Hallucinations typically manifest as misidentified objects, fabricated attributes, or omitted visual information, reducing the factual alignment between image and text.
@@ 6 Conclusion
Acknowledgements
The work is partially supported by the NSF of the United States Grant CRII 2451683, an NVIDIA Academic Grants Program, University of California at Merced, and a UC Merced Faculty Research Award. The views and conclusions are those of the authors and should not reflect the official policy or position of the U.S. Government.
We also thank Prof. Kai-Wei Chang (Department of Computer Science, UCLA) for his insightful feedback on earlier drafts, including suggestions on clarifying terminology, adding appropriate cognitive-science references, discussing experimental settings, and presenting qualitative examples, as well as for pointing us to relevant related work. Any remaining errors are our own.
@@ Appendix F Efficiency Comparison
We report MRFD’s efficiency on the CHAIR benchmark with two backbones—InstructBLIP-7B and LLaVA-1.5-7B—under the same decoding setup (128 max tokens, RTX 3090 GPU). With multi-region (K=3K{=}3) analysis and fusion, MRFD increases latency by about ${\sim}2.84\times$ (from 3.59 s to 10.21 s) and peak GPU memory by ${\sim}1.11\times$ (from 15,238 MB to 16,932 MB) relative to Regular decoding on InstructBLIP-7B (Table 8). On LLaVA-1.5-7B, MRFD shows a ${\sim}3.30\times$ latency increase (3.44 s $\rightarrow$ 11.34 s) and a ${\sim}1.11\times$ memory increase (15,778 MB $\rightarrow$ 17,458 MB) over Regular (Table 9). Despite this overhead—and compared to lighter baselines such as VCD or the post-hoc Woodpecker—MRFD remains markedly more efficient than iterative/feedback-heavy methods (DeGF, OPERA, HALC), while achieving the best CHAIRs (lower is better): 38.2 on InstructBLIP-7B and 37.1 on LLaVA-1.5-7B, representing ${\sim}33\%$ and ${\sim}32.5\%$ reductions from their respective Regular baselines. Overall, MRFD offers a compelling balance between computational cost and hallucination reduction.
Method Avg. Lat. $\downarrow$ GPU Mem. $\downarrow$ CHAIRs $\downarrow$
Regular 3.59 s 15,238 MB 57.0
VCD 6.82 s 16,078 MB 60.4
OPERA 25.31 s 23,575 MB 62.2
Woodpecker 11.72 s 22,798 MB 60.8
HALC 21.84 s 22,689 MB 53.8
DeGF 14.69 s 18,726 MB 49.2
Ours (MRFD) 10.21 s 16,932 MB 38.2
Table 8: Efficiency comparison on InstructBLIP-7B. Lower is better for all metrics.
Method Avg. Lat. $\downarrow$ GPU Mem. $\downarrow$ CHAIRs $\downarrow$
Regular 3.44 s 15,778 MB 55.0
VCD 6.91 s 16,634 MB 54.4
OPERA 24.70 s 22,706 MB 52.6
Woodpecker 10.68 s 22,199 MB 57.6
HALC 22.61 s 23,084 MB 51.0
DeGF 13.89 s 19,119 MB 48.8
Ours (MRFD) 11.34 s 17,458 MB 37.1
Table 9: Efficiency comparison on LLaVA-1.5-7B. Lower is better for all metrics.